OpenStudy (anonymous):

Help please?! Medal + Fan Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - y = -11 2x + 3y = 5 -4x - 9y = -19 -10y = -30 4x + 3y = 5 2y = -6 7x - 3y = -11 9x = -6 12x - 3y = -33 14x = -28

OpenStudy (anonymous):

I'm really not sure what the answer is.

OpenStudy (asevilla5):

hmm... here is what I did (I used substitution method) 4x - y = -11 --> y = 4x + 11 2x + 3y = 5 = 2x + 3(4x + 11) = 5 distribute 2x + 12x + 33 = 5 add like terms and subtract both sides by 33 14x = -28 <-- (shortcut: which option has that?)

OpenStudy (anonymous):

D has 12x - 3y = -33 14x = -28 so D has the -28

OpenStudy (asevilla5):

isn't that the solution given at the bottom and the 4 above that are the options?

OpenStudy (anonymous):

this is the question: Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - y = -11 2x + 3y = 5 --------------- A) -4x - 9y = -19 -10y = -30 B) 4x + 3y = 5 2y = -6 C) 7x - 3y = -11 9x = -6 D) 12x - 3y = -33 14x = -28

OpenStudy (asevilla5):

oh okay thanks for clarifying that. :)

OpenStudy (anonymous):

sorry for confusing you

OpenStudy (anonymous):

it's no problem :)